# Mental Disorders Detection in the Era of Large Language Models

### v1

@@ Appendix C Full Results on D-all and A-all Datasets
| D-all | Gemma2 2B IT | 5-shot | 19.63 | 94.12 | 32.49 | 92.86 | 16.56 | 28.11 | 30.3 | 100.0 | ||||||||||
| Gemma2 2B IT | 0-shot | 42.33 | 97.18 | 58.97 | 92.86 | 21.67 | 35.14 | 47.05 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 31.9 | 98.11 | 48.15 | 96.43 | 19.57 | 32.53 | 40.34 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 76.69 | 93.28 | 84.18 | 67.86 | 33.33 | 44.71 | 64.44 | 100.0 | |||||||||||
| Qwen2 7B IT | 5-shot | 21.47 | 94.59 | 35.0 | 92.86 | 16.88 | 28.57 | 31.79 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 54.6 | 94.68 | 69.26 | 82.14 | 23.71 | 36.8 | 53.03 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 5.52 | 100.0 | 10.47 | 100.0 | 15.38 | 26.67 | 18.57 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 15.95 | 96.3 | 27.37 | 96.43 | 16.46 | 28.12 | 27.75 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 39.88 | 92.86 | 55.79 | 82.14 | 19.01 | 30.87 | 43.33 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 2.45 | 100.0 | 4.79 | 100.0 | 14.97 | 26.05 | 15.42 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 57.06 | 96.88 | 71.81 | 89.29 | 26.32 | 40.65 | 56.23 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 31.29 | 91.07 | 46.58 | 82.14 | 17.04 | 28.22 | 37.4 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 18.4 | 93.75 | 30.77 | 92.86 | 16.35 | 27.81 | 29.29 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 17.18 | 96.55 | 29.17 | 92.86 | 16.88 | 28.57 | 19.25 | 95.81 | |||||||||||
| A-all | Gemma2 2B IT | 5-shot | 20.69 | 72.0 | 32.14 | 90.67 | 49.64 | 64.15 | 48.15 | 100.0 | ||||||||||
| Gemma2 2B IT | 0-shot | 26.44 | 69.7 | 38.33 | 86.67 | 50.39 | 63.73 | 51.03 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 45.98 | 66.67 | 54.42 | 73.33 | 53.92 | 62.15 | 58.28 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 45.98 | 60.61 | 52.29 | 65.33 | 52.13 | 57.99 | 36.76 | 98.77 | |||||||||||
| Qwen2 7B IT | 5-shot | 37.93 | 63.46 | 47.48 | 74.67 | 50.91 | 60.54 | 54.01 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 67.82 | 60.82 | 64.13 | 49.33 | 56.92 | 52.86 | 58.49 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 14.94 | 68.42 | 24.53 | 92.0 | 48.25 | 63.3 | 43.92 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 33.33 | 63.04 | 43.61 | 77.33 | 50.0 | 60.73 | 52.17 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 78.16 | 54.4 | 64.15 | 24.0 | 48.65 | 32.14 | 48.15 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 21.84 | 63.33 | 32.48 | 85.33 | 48.48 | 61.84 | 47.16 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 13.79 | 70.59 | 23.08 | 93.33 | 48.28 | 63.64 | 43.36 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 33.33 | 59.18 | 42.65 | 73.33 | 48.67 | 58.51 | 50.58 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 42.53 | 56.92 | 48.68 | 62.67 | 48.96 | 54.97 | 34.55 | 99.38 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 16.09 | 60.87 | 25.45 | 82.67 | 48.44 | 61.08 | 28.85 | 93.21 |

### v3

@@ Appendix C Full Results on D-all and A-all Datasets
| D-all | Gemma2 2B IT | 5-shot | 94.25 0.07 | 20.12 0.25 | 33.16 0.34 | 16.65 0.04 | 92.86 0.00 | 28.23 0.06 | 30.70 0.20 | 100.00 0.00 | ||||||||||
| Gemma2 2B IT | 0-shot | 96.06 0.56 | 41.84 0.25 | 58.29 0.34 | 21.00 0.33 | 90.00 1.43 | 34.05 0.54 | 46.17 0.44 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 5-shot | 98.17 0.03 | 32.88 0.49 | 49.26 0.56 | 19.80 0.12 | 96.43 0.00 | 32.85 0.16 | 41.06 0.36 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 0-shot | 93.28 0.00 | 76.69 0.00 | 84.18 0.00 | 33.33 0.00 | 67.86 0.00 | 44.71 0.00 | 64.44 0.00 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 5-shot | 94.59 0.00 | 21.47 0.00 | 35.00 0.00 | 16.88 0.00 | 92.86 0.00 | 28.57 0.00 | 31.79 0.00 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 0-shot | 95.64 0.48 | 56.56 0.98 | 71.08 0.91 | 25.17 0.73 | 85.00 1.43 | 38.84 1.02 | 54.96 0.96 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 5-shot | 100.00 0.00 | 7.48 0.98 | 13.91 1.72 | 15.66 0.14 | 100.00 0.00 | 27.08 0.21 | 20.50 0.97 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 0-shot | 96.18 0.06 | 15.46 0.25 | 26.64 0.37 | 16.38 0.04 | 96.43 0.00 | 28.01 0.06 | 27.32 0.21 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 93.02 0.08 | 40.86 0.49 | 56.78 0.49 | 19.26 0.13 | 82.14 0.00 | 31.21 0.17 | 43.99 0.33 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 100.00 0.00 | 2.45 0.00 | 4.79 0.00 | 14.97 0.00 | 100.00 0.00 | 26.05 0.00 | 15.42 0.00 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 96.88 0.00 | 57.06 0.00 | 71.81 0.00 | 26.32 0.00 | 89.29 0.00 | 40.65 0.00 | 56.23 0.00 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 90.94 0.06 | 30.80 0.25 | 46.01 0.28 | 16.94 0.05 | 82.14 0.00 | 28.08 0.07 | 37.05 0.18 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 93.90 0.08 | 18.90 0.25 | 31.46 0.35 | 16.43 0.04 | 92.86 0.00 | 27.93 0.06 | 29.69 0.20 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 96.73 0.09 | 18.16 0.49 | 30.58 0.70 | 17.59 0.35 | 95.71 1.43 | 29.71 0.57 | 20.10 0.43 | 95.81 0.00 | |||||||||||
| A-all | Gemma2 2B IT | 5-shot | 69.78 1.11 | 20.69 0.00 | 31.92 0.11 | 49.63 0.00 | 89.60 0.53 | 63.88 0.14 | 35.14 6.50 | 99.51 0.25 | ||||||||||
| Gemma2 2B IT | 0-shot | 71.44 0.87 | 26.44 0.00 | 38.59 0.13 | 50.69 0.15 | 87.73 0.53 | 64.26 0.27 | 51.42 0.20 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 5-shot | 67.80 0.57 | 38.62 3.68 | 49.09 2.66 | 52.52 0.70 | 78.67 2.67 | 62.96 0.40 | 56.02 1.13 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 0-shot | 60.61 0.00 | 45.98 0.00 | 52.29 0.00 | 52.13 0.00 | 65.33 0.00 | 57.99 0.00 | 36.76 0.00 | 98.77 0.00 | |||||||||||
| Qwen2 7B IT | 5-shot | 67.93 2.23 | 34.25 1.84 | 45.47 1.01 | 51.52 0.30 | 81.07 3.20 | 62.98 1.22 | 54.22 0.11 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 0-shot | 60.01 0.41 | 69.66 0.92 | 64.47 0.17 | 56.72 0.10 | 46.13 1.60 | 50.87 0.99 | 57.67 0.41 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 5-shot | 69.34 0.46 | 17.70 1.38 | 28.18 1.83 | 48.79 0.27 | 90.93 0.53 | 63.50 0.10 | 45.84 0.96 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 0-shot | 61.30 0.87 | 32.41 0.46 | 42.41 0.60 | 49.31 0.34 | 76.27 0.53 | 59.90 0.42 | 51.15 0.51 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 56.94 1.27 | 68.05 5.06 | 61.86 1.14 | 51.63 1.49 | 40.00 8.00 | 44.69 6.27 | 53.28 2.56 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 63.04 0.15 | 20.00 0.92 | 30.36 1.06 | 48.22 0.13 | 86.40 0.53 | 61.89 0.03 | 46.12 0.52 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 62.12 4.24 | 22.07 4.14 | 32.16 4.54 | 48.08 0.10 | 83.73 4.80 | 61.04 1.30 | 46.60 1.62 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 59.18 0.00 | 33.33 0.00 | 42.65 0.00 | 48.67 0.00 | 73.33 0.00 | 58.51 0.00 | 50.58 0.00 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 51.38 2.77 | 24.14 9.20 | 32.22 8.23 | 46.04 1.46 | 74.40 5.87 | 56.71 0.87 | 41.01 3.23 | 99.88 0.25 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 60.17 0.35 | 17.01 0.46 | 26.52 0.53 | 48.11 0.16 | 81.60 0.53 | 60.53 0.28 | 29.02 0.09 | 93.70 0.25 |
Appendix D Psychologists Evaluation Setup
During the LLM generation interpretation, psychologists used the following rating scale for explanations, given by LLM: 1 - completely erroneous, 2 - mostly erroneous, 3 - partially erroneous, partially correct, 4 - mostly correct, 5 - completely correct. On this scale, the psychologists rated LLM explanations as 2.84 out of 5 with Fleiss’ , which shows fair inter-annotator agreement. To better categorize errors in LLM explanations, psychologists also provided a list of the most common errors in each generation with detailed explanations of the error type. This list is provided in Table˜17.
| Error name | Description | Example |
| 1. Tautology | The final part of the explanation repeats the initial part in other words without any proof. | This text indicates that the author has depression, as he describes his thoughts and feelings, which are characteristic of depressive disorders. |
| 2. Groundless generalization | The patient’s experience from the text is defined as a sign of depression, while this experience on its own, without the context, is not specific to depression. | The desire to return to the lost state of happiness and happy life, which is also a sign of depression. |
| 3. False conclusion | The false inference is derived from the text statement. | The author also mentions that his parents, who he considers to be positive, were unable to correct his behavior. In the original text there is no signs of parents intention to correct authors behaviour. |
| 4. Confabulation | The explanation contains evidence, which is not mentioned in the context. | The text describes a deep dissatisfaction with the world, people and their actions included in the text, and also expresses a desire to be happy not in the text and enjoy the little things not in the text. |
| 5. Distortion of medical understanding of depression | Misconception about depression. | She also expresses a desire to … plan long-term plans, which also indicates the presence of a depressive disorder. The long-term planning cannot be considered as a sign of depression. |
| 6. Incompleteness of selected signs of depression | Of the several significant signs of depression only one or two signs (mostly minor) are highlighted. | Thank you for this test, so that I could repeat all this to myself once again and think about the rope. This significant sign does not mentioned in the explanation. |
We also conducted an additional experiment with a clinically informed prompt to explore how the additional information from psychologists can affect the quality of explanations. To do so, we asked psychologists to select parts of the text that can serve as depression markers for texts that were used in a 5-shot prompt for the DE dataset. After we performed a 5-shot evaluation of the best model with a modified 5-shot prompt, which now includes depression markers along with the target label. The results are presented in Table˜18.
| Corpus | Model | Mode |
F1-healthy |
F1-pathology | F1-macro | ||||||||
| DE | 0-shot | Vikhr 7B IT 5.4 | 71.43 0.00 | 5.56 0.00 | 10.31 0.00 | 19.05 0.00 | 90.91 0.00 | 31.50 0.00 | 20.90 0.00 | ||||||||
| 5-shot | Vikhr 7B IT 5.4 | 87.06 0.00 | 82.22 0.00 | 84.57 0.00 | 40.74 0.00 | 50.00 0.00 | 44.90 0.00 | 64.73 0.00 | |||||||||
| 5-shot clinically informed | Vikhr 7B IT 5.4 | 87.64 0.00 | 86.67 0.00 | 87.15 0.00 | 47.83 0.00 | 50.00 0.00 | 48.89 0.00 | 68.02 0.00 |
The adapted prompt leads to a slightly better F1-macro score; however, such a prompt requires qualified psychologists to label several texts for each investigated dataset. These results provide an interesting direction for future work and show that the overall quality of LLMs in mental disorders detection tasks can be improved with the help of psychologists.
To further investigate the quality of model explanations with clinically informed prompts, we asked psychologists to conduct the evaluation of results as in Section˜5.4. During evaluation, the psychologists set the average score to 2.47 out of 5, which is lower than for prompting without a clinically informed prompt. Due to the significant difference in the explanations of LLM, the mistakes were categorized into other groups, namely: (1) confabulation, (2) undifferentiation (inability to select symptomatic parts of the text), and (3) incompleteness of explanation. With a clinically informed prompt, each of the explanations contains at least one of the mentioned mistakes. The most common types of mistakes are confabulation and undifferentiation, which are found in 47% of the explanations. The incompleteness of the explanation occurs in 40% of the explanations. Overall, these results show that even with the clinically informed prompt, modern LLMs are unable to generate explanations that meet the requirements of clinicians.
Appendix E Feature Importance Ablation
To further deepen the explanation analysis, we conducted feature ablation for the best LLM on the DE dataset. For this purpose, we extracted the top-3 words by their importance on target label generation (and on full explanation generation) for each text in the test set. The words were extracted with the Feature Ablation method from the Captum framework Kokhlikyan et al. (2020), which calculates feature ablation based on differences in predictions with and without features. The list of unique words with the biggest mean feature importance is shown at Figure˜1 (importance for target label generation) and Figure˜2 (importance for full explanation generation). For both target class and explanation generation, the most important features in texts with predicted pathology class contain a significant amount of words with negative meanings (such as "disappointing", "negative", "bitter", "angry", "disgusting", "darkness", etc.), as well as words with direct pathology description ("depression", "fear"). The detailed analysis from the trained clinicians reveals that most of the features with high importance for depression class are connected with fear, suffering, and unhealthy conditions. On the other hand, the most important features of a healthy class contain mostly positive semantics, such as "humanity", "unselfishness", and "kind".
Appendix F Disaggregated Results for DSM Dataset
The scores for the Depression-Social Media (DSM) dataset were aggregated based on the obtained BDI questionnaire scores. To further investigate how these results vary between subgroups with various scores within one group (inside sub-splits of healthy or pathology class), we provided Tables˜19 and 20 with disaggregated scores. As one can see, for the healthy class, the F1-healthy scores are mostly consistent for all models, with the exception of RuBioRoBERTa, which has a significantly lower F1-healthy score for the split with BDI Score from 3 to 6. For pathology class, F1-pathology scores vary more significantly, which can be described with a bigger variability of initial scores, especially for the split with scores from 34 to 63.
| Mode | Model | BDI Score 0-3 | BDI Score 3-6 | BDI Score 7-10 | ||||||||||||||||
F1-healthy |
F1-healthy |
F1-healthy | ||||||||||||
| 5-shot | SaigaLlama3 8B | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 100.00 0.00 | 6.25 0.00 | 11.76 0.00 | ||||||||||
| 5-shot | Vikhr 7B IT 0.4 | 100.00 0.00 | 83.33 0.00 | 90.91 0.00 | 100.00 0.00 | 80.00 0.00 | 88.89 0.00 | 100.00 0.00 | 81.25 0.00 | 89.66 0.00 | ||||||||||
| 5-shot MMLU | SaigaLlama3 8B | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 100.00 0.00 | 6.25 0.00 | 11.76 0.00 | ||||||||||
| 5-shot MMLU | Vikhr 7B IT 5.4 | 100.00 0.00 | 83.33 0.00 | 90.91 0.00 | 100.00 0.00 | 80.00 0.00 | 88.89 0.00 | 100.00 0.00 | 81.25 0.00 | 89.66 0.00 | ||||||||||
| LoRA | Qwen2 7B IT | 100.00 0.00 | 80.56 11.45 | 88.79 7.00 | 100.00 0.00 | 83.33 21.34 | 89.15 15.15 | 100.00 0.00 | 65.62 10.05 | 78.80 7.34 | ||||||||||
| SFT | RuBioRoBERTa | 100.00 0.00 | 86.11 11.45 | 92.12 6.78 | 100.00 0.00 | 50.00 10.00 | 66.07 8.93 | 100.00 0.00 | 58.33 4.66 | 73.57 3.79 | ||||||||||
| Mode | Model | BDI Score 30-34 | BDI Score 34-63 | ||||||||||
F1-pathology |
F1-pathology | ||||||||
| 5-shot | SaigaLlama3 8B | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | ||||||
| 5-shot | Vikhr 7B IT 0.4 | 100.00 0.00 | 52.00 4.00 | 68.33 3.33 | 100.00 0.00 | 37.50 0.00 | 54.55 0.00 | ||||||
| 5-shot MMLU | SaigaLlama3 8B | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | ||||||
| 5-shot MMLU | Vikhr 7B IT 5.4 | 100.00 0.00 | 50.00 0.00 | 66.67 0.00 | 100.00 0.00 | 37.50 0.00 | 54.55 0.00 | ||||||
| LoRA | Qwen2 7B IT | 100.00 0.00 | 58.33 14.62 | 72.59 11.90 | 100.00 0.00 | 41.67 11.79 | 57.87 11.50 | ||||||
| SFT | RuBioRoBERTa | 100.00 0.00 | 55.00 5.00 | 70.83 4.17 | 100.00 0.00 | 27.08 4.66 | 42.42 5.42 | ||||||
Appendix G Textual Examples
For a better understanding of the investigated task, we presented several samples from the DE dataset - one per class. The samples were paraphrased and anonymized, and then translated into English. The examples are presented in Table˜21.
| Label | Text (In English) |
| Pathology | Hello! My name is NAME, I am already AGE years old. I live with my family: my husband and our wonderful daughter. My profession is a midwife in a women’s clinic, which I truly love. The team in our department is small, but we are all very close to each other. If any problems arise, we solve them very quickly. Until recently, I was an active person, with a large circle of friends who delighted me with various entertainments: going to the theater, watching movies, walking around the city or just meeting for a cup of coffee. Communicating with them brought me great pleasure. However, recently I began to distance myself from everyone, both at work and in my personal environment. This was the result of my fear, which turned out to be much more powerful than I expected. Fear takes away not only emotional strength, but also physical. At present, it seems to me that it is better to remain alone, avoiding meetings with other people. Although this makes my life less bright, I cannot cope with this condition yet. Also, I used to be into bead embroidery, but now I can’t even bring myself to do a simple task. Luckily, we have our cute house cat, who helps me cope with stress and is a great antidepressant. This is my current lifestyle. |
| Healthy | When I was a child, I accidentally found out that there were people who didn’t like me. They were my classmates, and I was very upset when I tried to be friends with them, but my attempt only increased the hostility. I tried to attract their attention with gifts, invitations to visit and other ways, but each new attempt ended in failure and even more trouble. Then I realized that no matter how you behave, there will always be people who don’t like you, sometimes for no apparent reason. But it is important to understand that this is normal and you shouldn’t try to live up to the expectations of others. You should value yourself for who you are. Many years have passed since then, and now I have many friends and acquaintances, some leave, and new ones come. Each person is unique and beautiful in their own way. I especially liked the expression: “Each person is a small cosmos.” It is really profound. Inside each person there is a whole universe consisting of his life experience, mistakes, disappointments, joys, defeats and small victories. If a person allows others to open up, it can be incredibly beautiful and exciting. In a world where we encounter many people every day, I want to believe that everyone respects each other, despite the fact that everyone may not like them. Even those we find unattractive may be dear to someone else. So there is no need to worry about not being loved, because we can find a common language and fill each other with love for life. We are all small universes living together in one big world, trying to get to know each other every day. |